Project Objective: This praject is to enhance the understanding of linear equation and system of linear equations. 1. Suppose that you usually consume some ice creams monthly. The quantity demanded (In economics, the quantity demanded of a goods or service is the total amount that buyers are willing to and able to purchase.) is closed related to many factors. Please list some of them you can think of and explain how they affect the quantity demanded. 2. One of the most important factors is the price of the goods or service. The following table called the demand schedule. describes how the number of ice creams you may consume each month depends on the price of it. Please complete the following table The price of ice-cream in dollar P) 0.00 0.50 1.00 1.50 2.00 2.50 3.00
1. Suppose that you usually consume some ice creams monthly. The quantity demanded (In economics, the quantity demanded of a gaads or service is the total amount that buyers are willing to and able to purchase.) is closed related to many factors. Please list some of them you can think of and explain how they affect the quantity demanded. 2. One of the most important factors is the price of the goods or service. The following table. called the demand schedule, describes how the number of ice creams you may consume each month depends on the price of it. Please complete the following table. in dollar (P) The quantity demanded of ice 19 16 13 For example, when the price of the ice cream is $0.00, you will consume 19 ice creams in one month. While the price is $0.50, you will consume 16 ice creams.
For example, when the price of the ice cream is $0.00, you will consume 19 ice creams in one month. While the price is $0.50, you will consume 16 ice creams Find the pattern of quantity and try to explain 3. Use x-axis to represent the quantity demanded, and y-axis to represent the price of ice crea graph about the price (P) and the quantity demanded (Q) and find the equatian the line In Economics, the curve is called the demand curve m. From the table above, draw the 4. Suppose that this month you get promoted and your salary increases. You will be able to consume one more ice cream under each price. Find the new demand curve when her salary increases. Furthermore, compare these two curves. where does the new demand curve shift? 5. Make a guess, if your salary decreases, how does the new demand curve shift? Why? 6. Similar as consumers, for the seller of ice creams Ben, the quantity supplied is also closely related to the price of the goods (In Economics the quantity supplied of a goods or service is the total amount that sellers are willing to and able to sell). The following table, called the supply schedule descibes how the amount of ice crearns that the
Answer #1

1) the various factors that affect the quantity demanded in the market for consumer   goods are price of the good, price of the related goods, availability of substitution , income , taste , season (specific to ice cream demand) and so on. As while purchasing we need to budget everything according to our income and a middle income person also looks for alternative available and make sure if he is purchasing thing at the price where the price is equal to the Marginal Utility derived from its use.that is MU = P.

2

PRICE $ THE QUANTITY DEMENDED FOR ICE CREAM
0.00 19
0.50 16
1.00 13
1.50 10  
2.00 7
2.50 4
3.00 1

TO find out this we need to make the demand linear equation for the same we make

Q=a+ bp where q= quantity demanded , a is the Q intercept   , Q is  the dependent variable , p is price which is independent variable . for framing equation we need quantity when price is Zero and ( is the inverse of the slope.why because we are measuring the chanfge in quantity resulting from change in price

which is \rightarrow \Delta q /  \Delta p and we know the slope is  \Delta y/  \Delta x and we have flipped  it as our x variable is q and y is price

to find it q2-q1/ p2-p1 = 13 - 16/ 1.00-0.50 = - 3 /0.50 =-6

so Q = a + bp that is a here will represent the quantity demanded at zero price which is 10

so Q = 19 - 6 (P)

Q= 19 -6 (1.50) = 19- 9 =10 AND SO ON

3)  

3.5 DEMAND CURVE 2.5 0.5 10 0 15 20

Quantity demanded   

demand curve equation is Q = 19 - 6 (P)

4) suppose there is an increase in the salary this will result in right ward shift in the demand curve and the new equation where the demand is 22 at zero $ we find the new demand equation Q = 20- 6 (p) on the same schedule that is provided by you

3.5 2.5 z original DEMAND CURVE 0.5 0 0 10 15 20 25